A wheel with a radius of 1.961m and a mass of 0.296kg rolls without sliding down a plane that tilts at an angle of 21.940 degrees. If the wheel is accelerated by a torque of 1.595 N x m, calculate the power if the wheel, which is shaped like a homogeneous cylinder, has a kinetic energy of 13.90J
1
Expert's answer
2021-05-31T17:01:59-0400


A wheel with a radius of 1.961m and a mass of 0.296kg rolls without sliding down a plane that tilts at an angle of 21.940 degrees. If the wheel is accelerated by a torque of 1.595 N x m, calculate the power if the wheel, which is shaped like a homogeneous cylinder, has a kinetic energy of 13.90J


"I\varepsilon=M=F_{fr}r"

where I is moment of inertia of the wheel,

"\varepsilon" is angular acceleration,

M is torque.

"ma=mgsin\alpha-F_{fr}"


For homogeneous cylinder:

"I=mr^2\/2"

For moving without sliding:

"a=\varepsilon r"

The kinetic energy:

"T=\frac{mv^2}{2}+\frac{I\omega^2}{2}"

"\omega=v/r"


Then:

"F_{fr}=M\/r"

"F=ma=mgsin\alpha-M/r"

"v=r\sqrt{\frac{2T}{mr^2+I}}=2\sqrt{\frac{T}{3m}}"


Power:

"P=Fv=2(mgsin\alpha-M/r)\sqrt{\frac{T}{3m}}"

"P=2(0.296\cdot9.8sin21.94\degree-1.595/1.961)\sqrt{\frac{13.9}{3\cdot0.296}}=2.14\ W"